Grade 12General Physics

experiment based on using vernier callipers

mukund kumar
15 Years agoGrade 12

1 Answer

Ashwin Sinha
15 Years ago

Dear Mukund Kumar,

Here is the experiment , using vernier callipers---------

Aim: To determine i) The volume of the given cylinder by measuring its length and diameter

                            ii ) The volume of the given sphere by measuring its diameter.

Apparatus : Vernier Calipers,Cylinder and sphere.

Description of Vernier Calipers: A Vernier calipers consists of mainly two parts i) A 2cm wide 15cm long rectangular metal strip .The left end bottom side of this strip consists of a fixed jaw 1 (A) and at the same end jaw 2(C) at the top of this strip. On the strip a scale (5)is graduated in Inches along the upper edge and another scale(4) is graduated in Centimeters along the lower edge. This is calledMain Scale ‘S’ .

Vernier Calipers

ii) A metal frame V called vernier slides over the Main Scale ‘S’ . At the bottom of this frame V a button 8(P) is attached,which helps to fix this vernier at any desired place on the main scale.This verier frame consists of jaw1 (B) at the bottom and a jaw 2(D) at the top .Two scales are graduated on this frame corresponding to two scales on the Main Scale ‘S’. The two scales 6 and 7 on the vernier are called Vernier scale.Vernier scale consists of equal number of divisions. When we move vernier frame over the main scale, a thin strip (3) will be projected out.The projection will be exactly equal to the distance between Jaws 1(AB) i.e the thickness of the object between jaws.

The lower jaws 1,1(AB) are used to measure the thickness or external diameter of the tubes,cylinders or spheres.

The upper jaws 2,2 (CD) are used to measure the inner diameters of hallow bodies like tubes or holes.

The thin strip ( 3) is used to measure the depth of the objects like test tubes.

Theory :  Principle of vernier calipers – N divisions on the vernier scale is equal to (N-1) divisions on the main scale.

N V.S.D = (N-1) M.S.D

1 V.S.D = \frac{(N-1)}{N} M.S.D

Least count (L.C) of vernier calipers : Minimum length or thickness measurable with the vernier calipers is called its least count.

Least count (L.C) = 1 M.S.D – 1 V.S.D

L.C = 1 M.S.D – \frac{(N-1)}{N} M.S.D

L.C = 1 M.S.D [ 1-\frac{(N-1)}{N}]

L.C = \frac{1 M.S.D}{N}  = \frac{S}{N}

Where S is the value of  one Main scale division and N is the number of equal divisions on the vernier scale.

Procedure : First we have to determine the least count count of the given vernier calipers.

To determine the volume of the  cylinder we have to determine a)the length of the cylinder and b) radius of the cylinder and substituting these values in the equation for the volume of the cylinder we can calculate it.

a) To determine the length of the cylinder : Given cylinder is held gently between jaws 1,1 of the vernier calipers.The reading on the main scale just before the zero of the vernier is noted.This is called Main scale reading (M.S.R).The number of division (n) on the vernier which coincides perfectly with any one of the main scale divisions is noted.This is called vernier coincidence (V.C).The vernier coincidence (V.C=n) is multiplied by least count to get the fraction of a main scale division.This is added to the main scale reading (M.S.R) to total reading or total length of the cylinder.

Total reading = M.S.R + (V.C\times L.C)

Take the readings,keeping the cylinders between jaws 1,1 at different positions.Post the values of M.S.R and vernier coincidence (n) in the table.Take at least 5 readings, get the average of these 5 readings which is mean length()of the cylinder.

b)To determine the diameter of the cylinder : Place the cylinder diametrically between the jaws 1,1 of the vernier calipers, as in the above case post the values of M.S.R and vernier coincidence (n) in the table. Take at least 5 readings, calculate the average of these readings which gives the mean diameter ( d=2r ) of the cylinder.

c) To determine the volume of the cylinder :Substituting the values of mean length () of the cylinder and mean diameter ( r) of the cylinder which is already determined, in the formula V = \pi r^2 l cm^3.

d)To determine the diameter of the sphere : The given sphere is held firmly between jaws of the vernier calipers, in such a way the points where the jaws are in contact with sphere should be the two extremes of the chord of the cylinder.Post the values of the M.S.R and vernier coincidence (n) in the table . Take at least 5 readings, calculate the average of these readings which gives the mean diameter (d=2 r ) of the sphere.

e)To determine the volume of the sphere :Calculate the radius of the sphere  r = d/2 .Substitute the value of  mean radius (r)  in the formula of the volume of the sphere V = \frac{4}{3} \pi r^3 cm^3 .

Precautions to be taken while doing the experiment : 1) Take the M.S.R  and vernier coincide every time without parallax error. 2)Record all the reading in same system preferably in C.G.S system. 3) Do not apply excess pressure on the body held between the jaws. 4) Check for the ZERO error.When the two jaws of the vernier are in contact,if the zero division of the main scale coincides with the zero of the vernier scale no ZERO error will be there.If not ZERO error will be there, apply correction.

Observations: 

I) Least count of Vernier calipers :

a) Value of 1 Main scale division  = 1 M.S.D = S = ……..  cm,

b) Number of divisions on the vernier scale   N= ……… cm,

Least count              L.C = \frac{S}{N}   = ………. cm.

II) Volume of the Cylinder :

a)Length of the cylinder :

S.No M.S.R                acm Vernier Coincidence   (n) Fraction   b=n*L.C Total Reading (a+b) cm
1.        
2.        
3.        
4.        
5.        

Average length of the cylinder l = ……… cm.

b) Diameter of the cylinder :

S.No M.S.R                acm Vernier Coincidence   (n) Fraction   b=n*L.C Total Reading (a+b) cm
1.        
2.        
3.        
4.        
5.        

Average diameter of the cylinder  d = 2r = …………. cm,

Mean radius of the cylinder           r = d/2 = ………….. cm,

Volume of the cylinder                   V = \pi r^2 l cm^3.

III ) Volume of the sphere :

a) Diameter of the sphere :

S.No M.S.R                acm Vernier Coincidence   (n) Fraction   b=n*L.C Total Reading (a+b) cm
1.        
2.        
3.        
4.        
5.        

Average diameter of the sphere  d = 2r = …………. cm,

Mean radius of the sphere           r = d/2 = ………….. cm,

Volume of the sphere                   V = \frac{4}{3} \pi r^3 cm^3.

Result : 1. Volume of the given cylinder  V= ……….. cm^3

2.Volume of the given sphere      V= ……….. cm^3.

 

 

    Hope it helps you dear........

 

                                              All The Best..........

Plz. Plz. Plz. don't forget to approve my answer by clicking 'Yes' given below, if you loved it...... Plz. don't forget to do it.... Plz........WinkSmileCool